If A and B are the zeroes of the polynomial 3x2 + 5x - 2, then form a quadratic polynomial whose zeroes are 2A and 2B

Grade:9

1 Answers

Ñïmísh Ñâñdúrkär
8 years ago
ANS is 3x2+10x-8=0
First find out roots(zeroes) of the given equation
We have, A+B=-5/3 and AB= -2/3
Now find out the values of A and B, by either substituting A in terms of B OR B in terms of A
The value of A comes out to be -2 and B comes out to be 1/3. Now 2A becomes -4 and 2B becomes 2/3.
substitute them in x2-(A+B)x+(AB)=0...THATS IT..!!